# Tandem Training for Language Models

### v1

@@ 2 Related work
Scalable oversight. There has been a recent push to enable scalable oversight, where strong models can be overseen by weaker agents, e.g., via decomposition (Christiano et al., 2018) or debate (Irving et al., 2018). Complementary lines replace scarce human labels with AI feedback or process-level signals, showing that supervising intermediate reasoning can improve correctness and auditability at scale (Bai et al., 2022; Paul et al., 2023; Lightman et al., 2023). Research on weak-to-strong generalization (Burns et al., 2024) demonstrates that training with weak supervision can nevertheless surface strong capabilities, formalizing conditions under which a weaker teacher or signal suffices. Our tandem training approach operationalizes scalable oversight inside the solution trajectory: instead of relying on external judges, decompositions, or teacher labels, we impose random handoffs to a weaker collaborator during RL and reward success only when the strong model’s reasoning is continuable by weaker agents, incentivizing intelligible solutions that are more amenable to oversight.
@@ 3 Method: tandem training
As laid out in Appendix A, tandem training can be viewed as a form of regularization. Akin to other regularization methods, noise injection (via handoffs to the junior model) is performed only during training; at test time, the tandem-trained senior model generates solutions on its own.
@@ 4.2 Models
Base models. Here we prompt the chat version of Llama-2-7b [^5] https://hf.co/meta-llama/Llama-2-7b-chat-hf in one of five languages $L$ (English, German, French, Bulgarian, Serbian), using a system prompt and two in-context question–answer demonstrations to explain the task and the expected output format and language (see Appendix B), followed by the GSM8K question. The entire prompt is provided in language $L$ (questions were translated ahead of time using GPT-4.1-mini), which effectively switches the model’s default language from English to $L$. As the base models were not specifically tuned for GSM8K, they do not produce $\ll\gg$ jargon. GSM8K test accuracy ranges between 12% (Serbian) and 24% (English).
@@ 4.3 Training and testing
When computing the log-likelihood of a rollout, one could also choose to mask tokens produced by the junior, as they were not produced by the (senior) model being trained, but for simplicity we did not perform masking in the experiments reported here. Turn-taking in tandem rollouts was carried out uniformly at random at the word level (cf. Sec. 3). The senior model was tuned using low-rank adapters on all linear layers (Hu et al., 2022). For hyperparameters, see Appendix C.
We emphasize that our goal is not to develop new RL algorithms, but to showcase the novel tandem training paradigm, which can, in principle, be used with any RL algorithm. Hence, it is advantageous if tandem training works even with a simple RL algorithm, such as the one presented here.
@@ 5 Results
Since the two-fold goal of tandem training is for the trained senior model to stop using jargon while keeping task accuracy high, we track accuracy as well as jargon over the course of training. We do so by storing checkpoints after every 10 gradient updates and using them to generate senior answers for the GSM8K test set. Accuracy is measured as the fraction of answers that provide the correct numerical solution. Jargon is measured in two ways: notational jargon is the fraction of answers that contain $\ll$ or $\gg$, whereas linguistic jargon is the probability of the senior’s input language $L_{\text{sen}}$ in its generated output, according to a language identification method based on a fastText model (see Appendix D). Next, we discuss results for each of the three settings described in Sec. 4.3.
Skill disparity. The leftmost column of Fig. 2 shows results for the skill disparity setting, where, during tandem training, an (English) GSM8K-specialist senior is paired with a base-Llama-2 junior prompted in English. We observe that, whereas before tandem training the senior uses notational jargon ($\ll\gg$) in nearly every answer (99%), it entirely stops doing so (0%) within 20 gradient updates. [^6] To confirm that jargon indeed disappears due to tandem training, rather than simply due to RL, we isolated the effects of tandem training and RL in a supplementary experiment where we trained the senior using the same RL method, but on rollouts produced by the senior alone. In this setup, jargon remains at 99% for all checkpoints, confirming tandem training, rather than mere RL, as the cause for vanishing jargon. At this point, accuracy is 33%, slightly down from the 39% achieved by the senior before tandem training, and considerably higher than the 24% achieved by the frozen junior. After this point, accuracy first decreases and then remains stable and largely above junior level. We hypothesize that the decrease is due to a distribution shift over the course of training: easier problems are more likely to be solved correctly, and since our simple RL method discards unsuccessful rollouts, the senior adapts to easier problems over the course of training. However, since the test set still follows the overall difficulty distribution, test accuracy suffers. [^7] This hypothesis is supported by the supplementary experiment of RL without tandem rollouts (cf. footnote 6): here, too, test accuracy decreased with more training (from 40% to 34%), pointing to the RL method, rather than tandem rollouts, as the cause of accuracy deterioration. While mitigating this issue (e.g., by also learning from unsuccessful rollouts) is an interesting direction for future work, we emphasize the key point: despite jargon having entirely vanished after 20 gradient updates, accuracy still reaches 85% of the senior’s original performance and surpasses junior performance by 38%.
Figure 3: Results for language disparity setting, where tandem-trained senior (rows) and frozen junior (columns) are Llama-2 models prompted in different languages. Plots show senior’s language use when applying checkpoints from tandem training to GSM8K test data. (Senior and junior accuracy shown in Appendix F, Fig. 4.) Takeaway: Tandem training reduces senior jargon (use of senior’s input language) without driving accuracy below junior level.
@@
Language disparity. Finally, we consider the language disparity setting, where both the senior and junior are base-Llama-2 models, but prompted in different languages $L_{\text{sen}}$ and $L_{\text{jun}}$, respectively. Here, the senior’s input language $L_{\text{sen}}$ is considered jargon. As observed in Fig. 3, tandem training leads the senior to abandon its jargon fast and adopt the junior’s language instead, typically within 20 gradient updates. In terms of accuracy (Appendix F, Fig. 4), we find that, when the junior beats the senior before training, the senior catches up; when they do equally well before training, it remains this way; and when the senior beats the junior before training, senior accuracy tends to decrease, but never below junior accuracy. The tenor is again that tandem training makes jargon disappear without driving accuracy below junior level. Note in Fig. 3 that the switch from $L_{\text{sen}}$ to $L_{\text{jun}}$ is never direct, but always passes through an intermediate phase (around gradient update 10) where the senior outputs English—which is neither $L_{\text{sen}}$ nor $L_{\text{jun}}$. In other words, English appears as a transitory lingua franca. Inspection of tandem rollouts from training shows that, in early training, rollouts often begin with a mix of $L_{\text{sen}}$ and $L_{\text{jun}}$, which seems to confuse the models and lead them to switch to English amidst rollouts. As such English completions tend to be more successful, the trained senior adopts this behavior, but the frozen junior maintains its tendency to produce $L_{\text{jun}}$, especially early on in rollouts. This, in turn, lets the senior eventually abandon English in favor of $L_{\text{jun}}$.
@@ 6 Discussion
Learning signals. Our current sequence-level objective attributes rewards to entire rollouts, which is simple and effective but coarse. Three complementary refinements target sharper, lower-variance credit signals. First, masking junior-authored tokens during likelihood and policy-gradient computations would prevent spurious credit from leaking onto text that the trainable senior did not produce, aligning gradients with the senior’s actual contributions. Second, learning from negative examples would leverage failed rollouts as informative counterfactuals: contrastive or penalty terms could down-weight the specific trajectories and specific segments most predictive of handoff failure. Third, token- and span-level preference learning (Deng and Mineiro, 2024) can localize blame and credit to the precise tokens where handoffs break, rather than pooling signal uniformly across the trajectory. Our work shows that tandem training is effective even when combined with a simple RL method, and we expect the benefits to increase further with more sophisticated RL methods.
@@ 10 Reproducibility statement
Data, models, training and testing procedure, and evaluation methodology are described in the main text. Hyperparameters, prompts, and further method details are listed as appendices. Model artifacts used to instantiate seniors and juniors are referenced in the main text. Upon publication, we will release all code, models, and data required to reproduce our results.
@@ Acknowledgments
This work was performed when Robert West and Ashton Anderson spent their sabbaticals as Visiting Researchers with the AI Frontiers group at Microsoft Research. Thanks to all our Microsoft colleagues for their thoughtful feedback on this work (in particular Saleema Amershi, Siddhartha Sen, and Tim Davidson) and for their warm welcome in Redmond and New York!
Appendix A Tandem training as regularization
Tandem training can be viewed as a regularization method: by injecting noise into the training process, it encourages simpler, more generalizable behavior. The approach is especially similar in spirit to dropout in neural-network training Srivastava et al. (2014), where neurons of the network being trained are randomly “muted” so the network learns not to over-rely on specific activation patterns. Similarly, in tandem training, the senior model is randomly “muted” (and replaced by the “noisier” junior) so senior learns not to over-rely on specific reasoning and speaking patterns. As in other regularization methods, including dropout, noise injection is switched off at test time.
Appendix B Llama-2 prompts
@@ Appendix C Hyperparameters
Appendix D Language identification
@@
Appendix E AI assistants statement
@@
Appendix F Supplemental results
Fig. 4 plots the senior and junior accuracy for the language disparity setting.

### v2

@@ 2 Related work
Scalable oversight. There has been a recent push to enable scalable oversight, where strong models can be overseen by weaker agents, e.g., via decomposition (Christiano et al., 2018) or debate (Irving et al., 2018). Complementary lines replace scarce human labels with AI feedback or process-level signals, showing that supervising intermediate reasoning can improve correctness and auditability at scale (Bai et al., 2022; Paul et al., 2023; Lightman and others, 2023). Research on weak-to-strong generalization (Burns et al., 2024) demonstrates that training with weak supervision can nevertheless surface strong capabilities, formalizing conditions under which a weaker teacher or signal suffices. Our tandem training approach operationalizes scalable oversight inside the solution trajectory: instead of relying on external judges, decompositions, or teacher labels, we impose random handoffs to a weaker collaborator during RL and reward success only when the strong model’s reasoning is continuable by weaker agents, incentivizing intelligible solutions that are more amenable to oversight.
@@ 3 Method: tandem training
Tandem training can be viewed as a form of regularization: by injecting noise into the training process, it encourages simpler, more generalizable behavior. The approach is especially similar in spirit to dropout in neural-network training Srivastava et al. (2014), where neurons of the network being trained are randomly “muted” so the network learns not to over-rely on specific activation patterns. Similarly, in tandem training, the senior model is randomly “muted” (and replaced by the “noisier” junior) so the senior learns not to over-rely on specific reasoning and speaking patterns. Akin to other regularization methods, including dropout, noise injection (here via handoffs to the junior model) is performed only during training; at test time, the tandem-trained senior model generates solutions on its own.
@@ 4.2 Models
Base models. Here we prompt the chat version of Llama-2-7b [^5] https://hf.co/meta-llama/Llama-2-7b-chat-hf in one of five languages $L$ (English, German, French, Bulgarian, Serbian), using a system prompt and two in-context question–answer demonstrations to explain the task and the expected output format and language (see Appendix A), followed by the GSM8K question. The entire prompt is provided in language $L$ (questions were translated ahead of time using GPT-4.1-mini), which effectively switches the model’s default language from English to $L$. As the base models were not specifically tuned for GSM8K, they do not produce $\ll\gg$ jargon. GSM8K test accuracy ranges between 12% (Serbian) and 24% (English).
@@ 4.3 Training and testing
When computing the log-likelihood of a rollout, one may also mask tokens produced by the junior, as they were not produced by the (senior) model being trained, but for simplicity we did not perform masking in the main experiments reported here. Turn-taking in tandem rollouts was carried out uniformly at random at the word level (cf. Sec. 3). The senior model was tuned using low-rank adapters on all linear layers (Hu et al., 2022). For hyperparameters, see Appendix B.
We emphasize that our goal is not to develop new RL algorithms, but to showcase the novel tandem training paradigm, which can, in principle, be used with any RL algorithm. Hence, it is advantageous if tandem training works even with a simple RL algorithm, such as the one presented here. (See Appendix E for an exploration of more complex variants, allowing for soft-masking of junior tokens and for not discarding incorrect rollouts.)
@@ 5 Results
Since the two-fold goal of tandem training is for the trained senior model to stop using jargon while keeping task accuracy high, we track accuracy as well as jargon over the course of training. We do so by storing checkpoints after every 10 gradient updates and using them to generate senior answers for the GSM8K test set. Accuracy is measured as the fraction of answers that provide the correct numerical solution. Jargon is measured in two ways: notational jargon is the fraction of answers that contain $\ll$ or $\gg$, whereas linguistic jargon is the probability of the senior’s input language $L_{\text{sen}}$ in its generated output, according to a language identification method based on a fastText model (see Appendix C). Next, we discuss results for each of the three settings described in Sec. 4.3.
Skill disparity. The leftmost column of Fig. 2 shows results for the skill disparity setting, where, during tandem training, an (English) GSM8K-specialist senior is paired with a base-Llama-2 junior prompted in English. We observe that, whereas before tandem training the senior uses notational jargon ($\ll\gg$) in nearly every answer (99%), it entirely stops doing so (0%) within 20 gradient updates. [^6] To confirm that jargon indeed disappears due to tandem training, rather than simply due to RL, we isolated the effects of tandem training and RL in a supplementary experiment where we trained the senior using the same RL method, but on rollouts produced by the senior alone. In this setup, jargon remains at 99% for all checkpoints, confirming tandem training, rather than mere RL, as the cause for vanishing jargon. At this point, accuracy is 33%, slightly down from the 39% achieved by the senior before tandem training, and considerably higher than the 24% achieved by the frozen junior. Thereafter, accuracy first decreases and then remains stable and largely above junior level. We discuss and mitigate the accuracy decrease in the final paragraph of this section.
Figure 3: Results for language disparity setting, where tandem-trained senior (rows) and frozen junior (columns) are Llama-2 models prompted in different languages. Plots show senior’s language use when applying checkpoints from tandem training to GSM8K test data. (Senior and junior accuracy shown in Fig. 4.) Takeaway: Tandem training reduces senior jargon (use of senior’s input language) without driving accuracy below junior level.
@@
Language disparity. Finally, we consider the language disparity setting, where both the senior and junior are base-Llama-2 models, but prompted in different languages $L_{\text{sen}}$ and $L_{\text{jun}}$, respectively. Here, the senior’s input language $L_{\text{sen}}$ is considered jargon. As observed in Fig. 3, tandem training leads the senior to abandon its jargon fast and adopt the junior’s language instead, typically within 20 gradient updates. In terms of accuracy (Fig. 4), we find that, when the junior beats the senior before training, the senior catches up; when they do equally well before training, it remains this way; and when the senior beats the junior before training, senior accuracy tends to decrease (see discussion in next paragraph), but never below junior accuracy. The tenor is again that tandem training makes jargon disappear without driving accuracy below junior level. Note in Fig. 3 that the switch from $L_{\text{sen}}$ to $L_{\text{jun}}$ is never direct, but always passes through an intermediate phase (around gradient update 10) where the senior outputs English—which is neither $L_{\text{sen}}$ nor $L_{\text{jun}}$. In other words, English appears as a transitory lingua franca. Inspection of tandem rollouts from training shows that, in early training, rollouts often begin with a mix of $L_{\text{sen}}$ and $L_{\text{jun}}$, which seems to confuse the models and lead them to switch to English amidst rollouts. As such English completions tend to be more successful, the trained senior adopts this behavior, but the frozen junior maintains its tendency to produce $L_{\text{jun}}$, especially early on in rollouts. This, in turn, lets the senior eventually abandon English in favor of $L_{\text{jun}}$.
Mitigating accuracy decrease. In the above results, senior accuracy tends to decrease towards junior accuracy over the course of tandem training. We hypothesize that the decrease is due to a gradual distribution shift: easier problems are more likely to be solved correctly, and since our simple RL method discards incorrect rollouts, the senior adapts to easier problems over the course of training. However, since the test set still follows the overall difficulty distribution, test accuracy suffers. This hypothesis is supported by a supplementary experiment where RL was performed with rollouts from the senior alone, rather than tandem rollouts (cf. footnote 6): here, too, test accuracy decreased with more training (from 40% to 34%), pointing to the RL method, rather than tandem rollouts, as the cause of accuracy deterioration. We thus experimented with two modifications to the RL method: (1) instead of discarding incorrect rollouts, include them in the log-likelihood objective with negative weight $c<0$; (2) instead of weighing senior and junior tokens equally in the objective, “soft-mask” junior tokens via a multiplicative factor $j<1$. Modification 1 enables learning from failure; modification 2 lets the senior focus more on its own than on the junior’s behavior. (Note that the simpler RL method used in the main experiments above corresponds to $c=0,j=1$.) In Appendix E we show that, by optimizing the hyperparameters $(c,j)$ on a validation set, the accuracy decrease is strongly mitigated, while jargon still vanishes entirely—the best of both worlds.
@@ 6 Discussion
Learning signals. Our current sequence-level objective attributes rewards to entire rollouts, which is simple and effective but coarse. Three complementary refinements target sharper, lower-variance credit signals. First, masking junior-authored tokens during likelihood and policy-gradient computations can prevent spurious credit from leaking onto text that the trainable senior did not produce, aligning gradients with the senior’s actual contributions. Second, learning from negative examples can leverage failed rollouts as informative counterfactuals: contrastive or penalty terms could down-weight the specific trajectories and specific segments most predictive of handoff failure. Third, token- and span-level preference learning (Deng and Mineiro, 2024) can localize blame and credit to the precise tokens where handoffs break, rather than pooling signal uniformly across the trajectory. Our work shows that tandem training is effective even when combined with a simple RL method, and we expect the benefits to increase further with more sophisticated RL methods. This expectation is supported by the initial experiments laid out at the end of Sec. 5 and in Appendix E, which show that the first two refinements listed above improve senior task accuracy.
@@ 10 Reproducibility statement
Experiments were run on machines with one or two Nvidia A100 GPUs with 80 GB of memory each. Data, models, training and testing procedure, and evaluation methodology are described in the main text. Hyperparameters, prompts, and further method details are listed as appendices. Model artifacts used to instantiate seniors and juniors are referenced in the main text. We release all code, models, and data required to reproduce our results.
@@ Acknowledgments
This work was performed when Robert West and Ashton Anderson spent their sabbaticals as Visiting Researchers with the AI Frontiers group at Microsoft Research. Thanks to all our Microsoft colleagues for their thoughtful feedback on this work (in particular Saleema Amershi, Siddhartha Sen, and Tim Davidson) and for their warm welcome in Redmond and New York! We also gratefully acknowledge the help of Raghav Singhal and Difan Jiao, who ran additional experiments during the review phase (Appendix E).
Appendix A Llama-2 prompts
@@ Appendix B Hyperparameters
Appendix C Language identification
@@
Appendix D AI assistants statement
@@
Appendix E Supplemental results
(a) Validation accuracy: skill disparity setting.
(b) Validation accuracy: skill and language disparity setting.
(c) Test results for skill disparity setting (top) and skill and language disparity setting (bottom).
Figure 5: Results for tuning the hyperparameters $(c,j)$ (defined in Appendix E) of the RL algorithm. (a) Validation accuracy of the tandem model for a range of hyperparameter choices, in the skill disparity setting. (b) Idem, in the skill and language disparity setting. (c) Test results achieved by the senior models selected based on validation accuracy. In comparison to Fig. 2 (columns 1 and 3), GSM8K accuracy remains higher, while jargon ($\ll\gg$ in the case of skill disparity, $\ll\gg$ and English in the case of skill and language disparity) still disappears.
While the main experiments of Sec. 5 used a variant of REINFORCE with binary rewards (see Sec. 4.3), we also explored the following generalization:
  1. 1.
    Instead of discarding incorrect rollouts, include them in the loss with negative weight $c<0$ (whereas correct rollouts are included with positive weight 1).
  2. 2.
    Instead of weighing senior and junior tokens equally in the loss, “soft-mask” junior tokens via a multiplicative factor $j<1$.
Modification 1 lets the senior learn from failure, whereas modification 2 lets the senior focus more on its own than on the junior’s behavior. The simpler RL method used in the main experiments corresponds to $c=0,j=1$.
We validated a range of $(c,j)$ combinations in order to investigate whether a judicious choice of these hyperparameters can mitigate or prevent the task-accuracy deterioration observed in Sec. 5 while still teaching the senior to avoid jargon. Importantly, we should not use any predefined notion of jargon in order to choose hyperparameters. One key reason for choosing GSM8K as a domain was precisely that it offers an exactly measurable notion of jargon ($\ll\gg$), which facilitates studying the effects of tandem training; but in real-world settings we generally do not know in what aspects—including jargon—senior and junior might differ.
We hence chose $(c,j)$ without referring to jargon and measured jargon only after the fact, to determine whether the corresponding senior model with the chosen hyperparameters meets the criteria of reducing jargon while maintaining high GSM8K task accuracy. Concretely, for each $(c,j)$, we tandem-trained a senior model for one epoch on a random sample containing 90% of the GSM8K training portion and validated the GSM8K task accuracy of the tandem (consisting of the trained senior and the frozen junior) on the remaining 10%. After training, we chose the $(c,j)$ for which the tandem had maximum validation accuracy on average over the last few training checkpoints. Here, the rationale is that, in order for a tandem—where senior and junior take random turns during rollouts—to have high GSM8K accuracy, the senior must have learned to adapt to the junior, which we take as a heuristic for choosing a suitable senior.
In these preliminary experiments, we restricted ourselves to the skill disparity setting and to the skill and language disparity setting with a French junior. The training runs used the same hyperparameter values listed in Appendix B, with the exception of effective batch size, which, per GPU capacity, was set to 160 and 152, respectively. The tandem model’s validation accuracies are shown in Fig. 5(a) and 5(b), respectively, leading us to choose (c=0.5,j=0.2)(c=-0.5,j=0.2) for the skill disparity setting and (c=0.3,j=0.35)(c=-0.3,j=0.35) for the skill and language disparity setting. Evaluating the chosen seniors (on their own, not in a tandem) on the test portion of GSM8K (Fig. 5(c)) shows that tandem training makes them stop using jargon ($\ll\gg$, and additionally English in the skill and language disparity setting) while keeping task accuracy considerably higher than in the main experiments: whereas in the main experiments, task accuracy in the skill disparity setting noticeably dropped from its initial 39% and stabilized around 30% (Fig. 2, first column), here it remained at around 40% (Fig. 5(c), top row); and whereas in the skill and language disparity setting with a French junior (Fig. 2, third column) it stabilized between 22% and 24%, here it remained above 30% (Fig. 5(c), bottom row).
Taken together, these supplemental results show that a slightly more sophisticated variant of the RL method of Sec. 4.3 more closely achieves the key promise of tandem training: to make the senior more intelligible without sacrificing performance. We thus see exploring even more advanced RL algorithms as a promising direction for future work.